What Celsius temperature, T2, is required to change the volume of the gas sample in Part A (T1 = 15 ∘C , V1= 1590 L ) to a volume of 3180 L ? Assume no change in pressure or the amount of gas in the balloon.
Answer #1

Given:

Ti = 15.0 oC

= (15.0+273) K

= 288 K

Vi = 1590 L

Vf = 3180 L

use:

Vi/Ti = Vf/Tf

1590.0 L / 288.0 K = 3180.0 L / Tf

Tf = 576 K

= (576 - 273) oC

= 303 oC

Answer: 303 oC